### v2

# MergeMix: A Unified Augmentation Paradigm for Visual and Multi-Modal Understanding

Abstract
Vision-language alignment in multi-modal large language models (MLLMs) relies on supervised fine-tuning (SFT) or reinforcement learning (RL). To align multi-modal large language models (MLLMs) in the post-training stage, supervised fine-tuning (SFT) is a stable choice but requires human annotations and lacks task generalizations, while Reinforcement Learning (RL) searches for better answers from reward signals but suffers from computational overhead and instability. To achieve balance among scalability, efficiency, and alignment generalizations, we propose MergeMix, a unified paradigm that bridges SFT and RL with an efficient Token Merge based Mixup augmentation. As for the Mixup policy, we generate contextual aligned mixed images with the corresponding labels according to the merged attention maps with cluster regions. Then, we enhance the preference-driven paradigm for MLLMs by building preference pairs with raw images and MergeMix-generated ones and optimizing the soft preference margin with the mixed SimPO loss. Extensive experiments demonstrate that MergeMix not only achieves dominant classification accuracy as an augmentation method but also improves generalization abilities and alignment of MLLMs, providing a new learning paradigm for preference alignment with training efficiency and stability.
1 Introduction
Multi-modal Large Language Models (MLLMs) (liu2024llava; bai2025qwen2.5; tong2024cambrian) have recently demonstrated remarkable capabilities in integrating visual and textual information, enabling a wide range of applications from visual question answering to multi-modal reasoning. Since MLLMs are typically pre-trained on massive web-scale datasets, forcing them to possess a wide range of knowledge and general reasoning capabilities, Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL)-based preference optimization (yang2025visionthink) have emerged as two primary paradigms for aligning MLLMs with human preferences and specific task requirements. However, SFT depends on high-quality instruction–response annotations and optimizes the likelihood of reference responses, which does not explicitly model relative preferences between outputs. RL-based methods such as RLHF are more preference-aware, but require an additional reward model that may introduce bias or be exploited by the reward signal.
Due to the shortcomings of data quality and training efficiency, some works (zhu2024seva; zhu2025nsft; luo2024probing; tan2025beyond; wang2024sima) try to build performance pairs for optimization. How to build the preference pair with control and high-quality data for model training is the remaining open question. For example, SeVa (zhu2024seva) proposed a preference optimization method by building a loser through some classic augmentation (i.e., RandomCrop). Then, select the different responses for optimizing the model by a DPO loss (rafailov2023dpo). However, these methods have two drawbacks: the augmentations are highly random, and the DPO loss cannot be related to the data, which means SeVa can only select useful training data. Those technical causes SeVa can not control the quality of the loser, which is harmful for some visual question answering tasks, and reduces the training data by selecting “hard negatives”. Hence, we investigate an interesting question: Is it necessary to propose novel techniques rather than some classical machine learning methods in the MLLM scenario?
In this paper, we revisit the mixup augmentation, which synthesizes mixed samples and corresponding labels with given mixing ratios. However, two main challenges arise as illustrated in Figure 1: (1) achieving an optimal trade-off between efficiency and performance of mixup augmentations that rely on saliency-based metrics, (2) extending the augmentation to MLLMs properly, from classical image corruptions to data-dependent samples. Motivated by these perspectives, we propose a novel training framework called MergeMix, which builds preference pairs for MLLM training through data augmentation methods and ranking loss, thereby bridging the gap between SFT and RL. Figure 2 shows the two scenarios of MergeMix. (a) We introduce MergeMix, a novel data augmentation that generates mixed samples through token merge techniques. A bipartite soft matching strategy captures similarity information that preserves contextual features, ensuring the mask retains useful information. Meanwhile, MergeMix links the merge ratio and mixing ratio, aligning mixed images with the corresponding labels, enabling precise mixing data generation. (b) We propose a preference-driven paradigm for MLLMs, where augmented samples are defined as non-preferred responses (Loser) and clean samples as preferred responses (Winner). This paradigm facilitates preference tuning via the mixed SimPO loss, and leverages the mixing ratio as the soft preference margin to enable adaptive optimization. Altogether, Figure 1 shows these contributions yield an efficient and effective training strategy that achieves stronger alignment with human preferences while preserving the stability and scalability of SFT. Since the optimization object has a direct relationship with augmentation, it obtains a more robust ability in calibration. Extensive experiments show that MergeMix, as a training-time augmentation paradigm, achieves competitive performance in both image classification and MLLM benchmark with favorable efficiency.
Our contributions can be summarized as:
  1. (a)
    We use token merging to obtain a local clustered attention map, enabling the generation of mixed images with cluster regions, a label re-scaling strategy aligned the mixed images with their corresponding labels, achieve well performance on both overhead and classification accuracy.
  2. (b)
    We enhance the preference tuning paradigm for supervised fine-tuning of MLLMs, where mixed images are treated as losers, the mixing ratio is used as a soft preference reward score, and optimize the model adaptively via the mixed SimPO loss.
  3. (c)
    We validate that our method achieves state-of-the-art on several image classification datasets and benchmarks, along with the advantages of our training paradigm on several MLLM benchmarks.
Figure 1: Efficiency and for MergeMix: (a) The training time vs. accuracy of mixup methods with the DeiT-Small model. (b) The image classification Top-1 accuracy vs. training epochs of different mixup methods on the CIFAR100 dataset with the DeiT-Tiny model. (c) The radar plot of the results on part VQA tasks by LLaVA-7B, LLaVA with SFT, and MergeMix.
2 Related Work
In this section, we introduce the existing mixup approaches for image classification and token compression approaches in multi-modal large language models for efficient training or inference.
Mixup Augmentations
The Mixup method mitigates model overfitting by generating augmented samples through mixing two different images within a mini-batch. Broadly, Mixup methods can be categorized into two types: Static, which relies on human priors or randomness, and Adaptive, a data-dependent type that leverages certain metrics to guide the mixing process. (i) Static: MixUp (zhang2017mixup) generates mixed samples via linear interpolation with $\lambda$. CutMix (yun2019cutmix) extends this idea from the global pixel level to a local patch level by constructing a mask of size proportional to $\lambda$ to mix images. ResizeMix (qin2020resizemix) ensures that features from at least one class are always preserved in the mixed sample by resizing the source image before mixing. Other methods, e.g., FMix (harris2020fmix), SmoothMix (lee2020smoothmix), GridMix (Baek2021gridmix), and StarMix (jin2024starlknet), focus on improving the mask to obtain more suitable mixed samples. (ii) Adaptive: SaliencyMix (uddinsaliencymix) employs a saliency extractor to identify informative patches in images for mixing. Attentive-CutMix (icassp2020attentive) and SuperMix (dabouei2021supermix) utilize a teacher model to guide mask generation. PuzzleMix (kim2020puzzle) and Co-Mix (kim2021co) generate appropriate masks based on gradient information obtained from a forward pass of the samples. AutoMix (liu2022automix) and AdAutoMix (qin2024adversarial) adopt an end-to-end bi-optimization paradigm to produce mixed samples. TransMix (chen2022transmix), SMMix (chen2023smmix), and MixPro (zhao2023mixpro) specifically enhance ViTs by computing attention scores from a forward pass to generate feature-aware masks and further refine the label ratio through attention scores. DiffuseMix (islam2024diffusemix) and GenMix (islam2024genmix) generate mixed samples by a diffusion model (zhu2025obs) for label preserving.
Token Compression in MLLMs
Token compression methods (bolya2023tome; tao2025dycoke; shao2025holitom; chen2025streamingtom; tao2025omnizip) mainly propose to merge or drop redundant tokens to achieve efficiency and acceleration. In MLLMs, images and texts will incur a significant number of tokens, which are often full of redundant information. Obvious researchers bring the token compression into MLLMs. Overall, we divide the methods that reduce tokens into 2 types, Reduce in Encoder and Reduce in Decoder. (i) Reduce in Encoder: MADTP (cao2024madtp) aims to achieve MLLM acceleration by purging visual tokens. LLaVA-PruMerge (shang2024llavapruerge) uses the attention of [CLS] token to select clustering centers and then merges the remaining tokens with lower attention through a KNN clustering and weighted clustering center updating mechanism. VisionZip (yang2025visionzip), instead, retains visual tokens with high attention scores and subsequently merges the remaining tokens through clustering. Others, such as TokenPacker (li2025tokenpacker), AVG-LLaVA (lan2025avg), MustDrop (liu2024mustdrop), and LLaVolta (chen2024LLaVolta), achieve acceleration by choosing a metric to sample TopK visual tokens. FastVLM (vasu2025fastvlm) proposes an Efficient Vision Encoder to reduce visual tokens. (ii) Reduce in Deocder: PyramidDrop (xing2024pyramiddrop) divides the token compression process in LLM into multiple stages, which employs a pyramidal token drop to avoid losing too much visual information in shallower layers. ATP-LLaVA (ye2025atp) proposes an Adaptive Token Pruning (ATP) module that reduces the number of tokens in the decoder layer. ZipVL (he2024zipvl) proposes a dynamic ratio allocation strategy via the importance token, adaptively determined based on the distribution of attention scores in a particular layer, rather than a fixed hyperparameter.
3 Preliminaries
Reformulation of Mixup Augmentation.
We define $\mathbb{X}$ to be the set of training samples and $\mathbb{Y}$ the set of ground truth of the corresponding labels. For each sample pair $(x,y)$, we randomly sample two pairs in $\mathbb{X}$ and $\mathbb{Y}$, with $\lambda$ in $\text{Beta}(\alpha,\alpha)$. The mixed images and labels are generated by applying the optimized mask $\mathcal{M}$ and ratio $\hat{\lambda}$, which come from a defined policy $\mathcal{P}(\cdot,\cdot,\cdot)$ according to Eq. (1):
$\mathcal{M},\ \hat{\lambda}$ $=\mathcal{P}\big(f_{\theta}(x_{i},x_{j}),\ (y_{i},y_{j}),\ \lambda\big),$ (1)
$\hat{x}$ $=\mathcal{M}\odot x_{i}+(1-\mathcal{M})\odot x_{j},$ (2)
$\hat{y}$ $=\hat{\lambda}*y_{i}+(1-\hat{\lambda})*y_{j},$
where the $\odot$ denotes element-wise multiplication. Policy $\mathcal{P}(\cdot,\cdot,\cdot)$ aims for the $\mathcal{M}$ to retain more features in the mixed sample. The $\hat{\lambda}$ keeps the initial sampling ratio when without optimization; otherwise, the $\lambda$ can be re-computed by some metrics (i.e., MergeMix uses the total mask values).
Preference Tuning for MLLMs.
Preference optimization methods aim to align LLMs and MLLMs with human feedback by contrasting preferred and dispreferred responses. A general preference loss can be abstractly defined as Eq. (3):
$\mathcal{L}_{\text{Pref}}=-\log\sigma\big(\pi_{\theta}(x,y^{+})-\pi_{\theta}(x,y^{-})\big),$ (3)
where $(x,y^{+})$ and $(x,y^{-})$ denote the preferred and dispreferred responses respectively, and $s_{\theta}$ is a scoring function that reflects model preference. Different approaches (e.g., PPO (schulman2017ppo), DPO (rafailov2023dpo)) instantiate $\pi_{\theta}$ in various ways, but all share the same principle.
Unlike RL-based approaches, which require training a separate reward model, DPO provides a simple and stable alternative by directly optimizing the policy model using preference pairs. Formally, the DPO loss is defined as Eq. (4):
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}{(x,y^{+},y^{-})\sim\mathcal{D}}\Big[\log\sigma\big(\beta\log\frac{\pi_{\theta}(y^{+}|x)}{\pi_{\text{ref}}(y^{+}|x)}-\beta\log\frac{\pi_{\theta}(y^{-}|x)}{\pi_{\text{ref}}(y^{-}|x)}\big)\Big],$ (4)
where $\pi_{\theta}$ denotes the policy model (in our case, an MLLM such as LLaVA-v1.5 (liu2024llava) or Qwen-VL (bai2025qwen2.5)), and $\pi_{\text{ref}}$ represents a frozen reference model used to preserve alignment with the original pre-trained distribution. $\sigma$ is the sigmoid function, and $\beta>0$ is a temperature-like scaling factor that controls the sharpness of preference separation. Intuitively, DPO encourages the policy to assign a higher likelihood to preferred responses ($y^{+}$) than to non-preferred ones ($y^{-}$), while maintaining proximity to the reference model.
Figure 2: The overall of the two scenarios of MergeMix: (a) MergeMix for Image Classification: The image is processed by the ToMe encoder, with Attention Score Recovery and TopK sampling to generate the corresponding class prediction. (b) MergeMix for MLLM: Preference pairs are encoded by the vision model with token merging, and the LLM decoder generates response text for the loser and winner, optimized via a ranking loss.
4 MergeMix Training Paradigm
In this section, we present the implementation of MergeMix, an augmentation approach via token merging for image mixing, not only for image classification, but also designed for multi-modal large language models. Figure 3 shows the overall pipeline of MergeMix, and we describe in two subsections in detail, which are from the input space to the loss objective for model training.
4.1 Image Mixing via Token Merge
In MergeMix, we leverage the relationship between the merge ratio and mixing ratio. The merge ratio measures the information of raw samples, while the mixing ratio balances the information between mixing samples, thereby enabling precise data generation of mixed inputs and labels. In this subsection, we first introduce MergeMix on the input space. Then we use the designed mixing policy $\mathcal{P}(\cdot,\cdot,\cdot)$ to obtain the mixed images $\hat{x}$ with the mask $\mathcal{M}$.
Image Policy with Token Merging.
[…]
$S,A_{K},Z_{K}=\text{ToMeAttention}(Z_{L},r),$ (5)
where $A_{K}$ denotes the attention map from the model, and $Z_{K}$ denotes the feature tokens for computing one-hot loss. $r$ denotes the number of merged tokens, which can reduce some high-similarity semantic tokens and retain a condensed token sequence. Also, based on Token Merge, we obtained a source map $S$ for their spatial relationships between the raw token sequence $Z_{L}$ and the $Z_{K}$.
Generating Mixing Mask with Source Matrix.
Since token merging aggregates non-similar tokens into compact representations $Z_{K}$, the resulting matrix preserves local feature structures more effectively. In contrast, the vanilla TopK selection adopts a greedy sampling strategy with linear complexity $\mathcal{O}(N)$, which discards low-ranked tokens directly and thus loses spatial relationships. Alternatively, the Bipartite Soft Matching (BSM) approach performs global pairwise matching with quadratic complexity, yielding a more balanced and globally optimal merging of tokens. To reconstruct the full-resolution attention map, we introduce a recovering function $\mathcal{R}_{K\to L}(\cdot,\cdot)$ that expands the merged attention map $A_{K}$ back to its original length $A_{L}$, which shows in Fig 5 and according to Eq. (6):
$\hat{A_{L}}=\mathcal{R}_{K\to L}(A_{K},S).$ (6)
Unlike discrete TopK sampling, our recovery mechanism propagates merged attention over the original token topology guided by similarity $S$, restoring richer spatial dependencies and contextual continuity, thus reducing information loss from hard selection. Based on the encoder with a token merge and attention recovery. We can generate the binary $\mathcal{M}$ according to Eq. (7):
$\mathcal{M}_{i}=\left\{\begin{matrix}1,\ &\text{if }i\in\text{TopK}(\hat{A}_{L},p),\\ 0,\ &\text{otherwise},\end{matrix}\right$. (7)
where $p$ denotes the selection number, $p=\lfloor\lambda*L\rfloor$, and $i$ denotes the index of sequence. Finally, we can mix the mini-batch and get the augmented data for $f_{\theta}(\cdot)$ training.
Figure 3: Overall illustration of MergeMix for MLLM. (a) MergeMix performs attention-based mask mixing guided by the ToMe Vision Encoder, recovering token attention scores and generating a mixed image through an augmenter. Specifically, Token Merging hierarchically merges visual tokens via Bipartite Soft Matching (BSM) to enhance efficiency, which is trained with both the SFT and ranking losses. (b) Case study of preference data generated by MergeMix with LLaVA-v1.5-7B.
4.2 A Unified Augmentation Paradigm: From Image Classification to MLLMs
In this subsection, we describe the loss function $\mathcal{L}$. For the classification task and visual understanding, our final loss $\mathcal{L}_{\text{Total}}$ combines two losses: the main loss (one-hot cross entropy loss $\mathcal{L}_{\text{CE}}$ and $\mathcal{L}_{\text{SFT}}$) and the reformulated loss (mixup cross entropy loss $\mathcal{L}_{\text{MCE}}$ and ranking loss $\mathcal{L}^{\text{Mix}}_{\text{SimPO}}$). Figure 3 shows the pipeline of MergeMix for MLLM in detail.
Re-scaling Policy for Mixing Ratio.
Under this optimization objective, the role of the mixing ratio $\lambda$ is to serve as a metric that quantifies the presence of feature information from the two samples. While this metric cannot directly reflect the true characteristics of the data, certain adaptive methods can constrain the model to generate mixed samples where the mixing ratio progressively approximates the target value (jin2024survey). Some works, like LUMix (sun2022lumix), DecoupleMix (liu2023decoupledmix), and SUMix (qin2024sumix), use a defined policy for some hand-crafted mixup methods, and find that it is more efficient than optimizing a better mask way.
[…]
$\hat{\lambda}\sim\mathcal{N}(\mu,\sigma),\quad\hat{\lambda}=\text{clip}\big(\frac{\hat{\lambda}-\min(\hat{\lambda})}{\max(\hat{\lambda})-\min(\hat{\lambda})+\tau},0,1\big),$ (8)
where the $\mathcal{N}(\cdot,\cdot)$ denotes a Gaussian function, $\mu$ and $\sigma$ represent the merged ratio and mixing ratio repetitively. $\tau$ as a hyperparameter, set to 1e-5. Then, we obtain the re-scaled mixing ratio $\hat{\lambda}$ with spatial and ToMe model inherent features to optimize the model when training. In total, the loss of mixup training as:
$\mathcal{L}_{\text{Total}}=\underbrace{\mathcal{L}_{\text{CE}}\big(f_{\theta}(\hat{x}),y_{i}\big)\ast\hat{\lambda}+\mathcal{L}_{\text{CE}}\big(f_{\theta}(\hat{x}),y_{j}\big)\ast(1-\hat{\lambda})}_{\text{mce loss}}+\underbrace{\mathcal{L}_{\text{CE}}\big(f_{\theta}(x),y\big)}_{\text{one-hot loss}}.$ (9)
Aggregating Mixing Ratio within Preference Loss.
LLaVA (liu2024llava) uses a standard conditional language modeling loss for SFT. In MLLMs, we are given an instruction–response pair $(x,y)$, where $x$ denotes the multi-modal data, inducing vision and text, and $y=(y_{1},y_{|y|})$ denotes the target response. The SFT loss is defined as:
$\mathcal{L}_{\text{SFT}}=-\mathbb{E}_{(x,y)\sim\mathcal{D}}\Bigg[\sum_{t=1}^{|y|}\log\pi_{\theta}\big(y_{t}\mid x,y_{<t}\big)\Bigg].$ (10)
This objective needs to maximize the likelihood of GT responses, aiming to align the data. In Section 3, we introduced the DPO loss, which can be decomposed into two components: the SFT part and the ranking optimization part. In our approach, we replace the ranking component with SimPO (meng2024simpo), where $y$ denotes the target sequence (response) and $|y|$ denotes its length. Furthermore, since $\lambda$ reflects information similarity between augmented and raw image (interpreted as “loser degree” in MLLMs), we link it to $\gamma\to 1-\hat{\lambda}$: larger $\lambda$ represent higher similarity and harder discrimination, reduces $\gamma$ to avoid over-optimization on trivial differences; smaller $\lambda$ represent greater dissimilarity and easier tasks increases $\gamma$ to strengthen constraints for clearer preference distinction. The mixed SimPO loss replacement is Eq. (11):
$)].\displaystyle\mathcal{L}^{\text{Mix}}_{\mathrm{SimPO}}=-\mathbb{E}_{(x,\hat{x},y)\sim\mathcal{D}}\Big[\log\sigma\big(\frac{\beta}{|y|}\log\pi_{\theta}(y\mid x)-\frac{\beta}{|y|}\log\pi_{\theta}(y\mid\hat{x})-(1-\hat{\lambda}$$\big)\Big$]. (11)
This reformulated loss strictness with sample difficulty, enabling more robust preference optimization. Finally, the total loss of our training paradigm is written as:
$\mathcal{L}_{\text{Total}}=\mathcal{L}_{\text{SFT}}+\mathcal{L}_{\text{SimPO}}^{\text{Mix}}.$ (12)
5 Experiments
5.1 State-of-the-art methods.
Image classification.
To evaluate the performance of MergeMix, we compared with some mainstream mixup methods, i.e. Mixup (zhang2017mixup), CutMix (yun2019cutmix), FMix (harris2020fmix), SmoothMix (lee2020smoothmix), GridMix (Baek2021gridmix), ResizeMix (qin2020resizemix), SaliencyMix (uddinsaliencymix), Attentive-CutMix (icassp2020attentive), PuzzleMix (kim2020puzzle), GuidedMixup (kang2023guidedmixup), AutoMix (liu2022automix) and AdAutoMix (qin2024adversarial). DeiT (icml2021deit), TransMix (chen2022transmix), SMMix (chen2023smmix), MixPro (zhao2023mixpro) and TdAttenMix (wang2025tdattenmix) for some ViT-based methods. The training configures about datasets and methods follows the open-source library OpenMixup (2022openmixup).
MLLMs.
To evaluate the training paradigm that we proposed, we compare with three different system-level baselines: (1) SFT with different training paradigms on LLaVA, including LLaVA-NeXT-7/13B (liu2024llavanext), SeVa-7B (zhu2024seva), SIMA (wang2024sima), and nSFT (zhu2025nsft). (2). Token reduction on LLaVA, including LLaVA-PruMerge+ (shang2024llavapruerge), VisionZip (yang2025visionzip), VisPrunner (zhang2024vispunner), VScan (zhang2025vscan), and LLaVA-Mini (zhang2025llava-mini). (3). RL training on Qwen2.5-VL-Instruction (bai2025qwen2.5), including VisionThink (yang2025visionthink).
For all classification results, we report the top-1 test accuracy in the last 10 training epochs for each trial. To facilitate comparison, we mark the best and second best results in bold and cyan. For the LLaVA benchmark, we use the LLaVA official code, and for the Qwen2.5-VL-Instruction benchmark. We use lmms-eval (zhang2024lmmsevalrealitycheckevaluation) for evaluation.
Table 1: Top-1 accuracy (%) of mixup methods training 200 epochs on CIFAR100 dataset with different model sizes, T/S/B/L denotes Tiny, Small, Base, and Large, respectively. The full results of training 600 epochs are in Table A1.
Method DeiT-T DeiT-S ViT-S ViT-B ViT-L
Vanilla 64.70 65.81 62.64 63.33 61.83
MixUp 69.47 69.98 68.67 69.66 67.90
CutMix 75.98 74.21 69.67 72.18 68.97
FMix 72.73 70.41 68.41 68.62 66.12
GridMix 71.54 68.86 70.15 66.63 63.20
ResizeMix 69.42 68.54 67.86 63.72 63.48
SaliencyMix 69.83 69.78 70.14 68.75 67.12
PuzzleMix 73.40 73.60 70.92 71.13 69.77
AutoMix 72.91 76.24 68.44 73.40 72.10
AdAutoMix 72.83 72.63 69.66 71.43 69.69
DeiT 74.01 75.92 72.96 72.15 69.23
TransMix 75.31 76.17 74.15 72.87 71.40
SMMix 73.84 74.09 73.50 70.87 71.38
MixPro 74.78 75.26 73.49 73.18 72.28
TdAttenMix 73.63 73.32 73.11 72.19 72.12
MergeMix 77.46 78.68 77.02 75.75 76.19
Table 2: Top-1 accuracy (%) of mixup methods on the Stanford-Cars dataset. Full results of the CUB200 and FGVC-Aircrafts dataset in Table A2.
Method $\alpha$ DeiT-S ViT-B
Vanilla - 86.77 91.31
MixUp 1.0 87.73 91.36
CutMix 0.2 88.37 91.53
SmoothMix 0.2 86.39 90.88
FMix 0.2 87.18 91.36
GridMix 0.2 87.58 91.31
ResizeMix 1.0 87.45 91.59
Attentive-CutMix 2.0 87.35 90.29
SaliencyMix 0.2 87.94 91.47
PuzzleMix 1.0 88.60 91.83
[…]
DeiT 0.2 88.72 92.17
TransMix 1.0 88.38 91.66
SMMix 1.0 88.76 91.93
MixPro 1.0 88.38 91.48
TdAttenMix 1.0 88.78 91.68
MergeMix 1.0 89.42 92.20
[…]
Method DeiT-Small
[…]
Vanilla 1375.80 4.24 75.66
MixUp 1374.54 4.24 77.80
CutMix 1374.61 4.24 80.13
DeiT 1374.20 4.24 79.80
TransMix 1375.17 4.24 80.44
SMMix 1373.93 4.24 79.36
MixPro 1373.62 4.24 79.33
MergeMix 1591.66 3.56 80.71
5.2 Datasets
In our paper, we mainly divided into 2 scenarios: Image Classification and MLLM Benchmark. The detailed information about datasets is described in Appendix B.1. For the image classification datasets, we choose 5 public classification datasets, including the small-scale dataset of CIFAR100 (krizhevsky2009learning), the large-scale dataset of ImageNet-1K (russakovsky2015imagenet), and the fine-grained datasets of CUB200 dataset (wah2011caltech), FGVC-Aircrafts dataset (maji2013fine), and Stanford-Cars dataset (jonathan2013cars). For the MLLM datasets, we choose 16 datasets, including visual question answering (VQAv2 (goyal2017vqav2), GQA (hudson2019gqa), VizWiz (gurari2018vizwiz), ScienceVQAI (lu2022scivqa), TextVQA (singh2019textvqa), MME-RealWorldQA (zhang2025realworldqa)), understanding (MME (Perception) (yin2023mme), MMBench (liu2025mmbench), $^{\text{CN}}, MMBench$$^{\text{CC}}$, POPE (F1 score) (li2023pope), SEEDI (li2023seed), MMStar (chen2024mmstar)), and reasoning (MMMU (yue2024mmmu), MMMU-Pro Standard (MMMU-Pros) (yue2024mmmu-pro), MathVista (lu2024mathvista)).
Table 4: Full system-level comparison results in LLaVA. Compared with their counterparts. AVG denotes the average of the nine benchmarks for comprehensive comparison, except for MME, underline denotes MME with the sum of Perception and Cognition. Tokeni denotes training with the token number. Full results in Table A10, Table A11 and Table A12.
Models Tokeni Image Question Answering Benchmarks AVG Gain
[…]
LLaVA Variants
LLaVA-7B Full 78.5 62.0 50.0 66.8 58.2 1510.7 64.3 58.3 85.87 66.19 65.57 -
LLaVA-NeXT-7B Full 81.8 64.2 57.6 70.1 64.9 1519.0 67.4 60.6 86.5 70.2 69.3 -
LLaVA-NeXT-13B Full 82.8 65.4 60.5 73.6 67.1 1575.0 70.0 64.4 86.2 71.9 71.3 -
SeVa-7B Full - 60.7 - 67.5 56.2 1450 65.6 59.2 86.7 65.8 - -
SIMA Full - 62.2 54.4 68.1 58.3 1507.7 64.9 59.0 86.5 65.9 - -
nSFT Full - 62.9 - 68.5 58.7 1531 67.1 61.0 86.8 66.2 - -
LLaVA with Token Compressions
LLaVA-PruMerge+ 144 76.8 - - 68.3 57.1 1462.4 64.9 - 84.0 - - -
VisionZip 192 77.4 60.1 - 68.2 57.8 1834.0 63.4 - 84.9 57.1 - -
VisPrunner 128 75.8 58.2 52.7 69.1 57.0 1461.4 62.7 57.3 84.6 - - -
VScan 192 77.8 60.6 50.4 68.6 57.7 1806.0 63.9 57.4 86.2 - - -
LLaVA-Mini 1 77.6 60.9 56.2 70.4 57.0 1466.0 65.6 - 84.4 58.5 - -
LLaVA with Augmentations & Ranking Loss
SFT Vision Full 79.32 62.98 47.45 70.05 57.17 1490.88 66.26 60.05 86.18 67.32 66.31 +0.74
+ MixUp Full 79.27 62.58 44.95 69.41 57.39 1483.20 65.72 58.24 86.27 66.73 65.62 +0.05
+ CutMix Full 79.18 62.40 45.04 70.60 57.06 1452.31 66.32 58.24 86.47 67.22 65.84 +0.27
+ ResizeMix Full 77.78 61.66 44.43 68.91 55.11 1436.09 63.91 55.41 86.01 63.91 64.13 -1.44
+ MergeMix Full 79.24 62.44 47.69 69.86 57.56 1479.97 66.58 60.65 86.10 67.47 66.40 +0.83
SFT Vision 288 78.6 62.47 48.15 69.51 56.41 1486.24 66.32 57.98 87.37 66.75 65.95 +0.38
+ MixUp 288 78.51 62.07 51.1 68.47 56.54 1459.06 65.63 59.53 86.86 66.06 66.08 +0.51
+ CutMix 288 78.58 62.39 50.53 70.2 55.95 1414.72 66.92 59.53 86.56 66.2 66.31 +0.74
+ ResizeMix 288 76.39 61.05 45.48 68.07 54.60 1447.35 63.31 51.97 86.57 62.54 63.33 -2.24
+ MergeMix 288 78.61 62.18 52.14 69.61 56.85 1453.97 66.58 59.02 86.47 66.63 66.45 +0.88
Table 5: Full system-level comparison results in Qwen2.5-VL-Instruction (Qwen2.5-VL-Ins). AVG denotes the average of the nine benchmarks for comprehensive comparison.
[…]
Qwen2.5-VL-Ins-7B 62.42 84.02 80.41 62.94 86.38 68.63 50.3 36.42 19.2 61.19 -
VisionThink-7B 61.00 82.73 81.01 64.5 87.65 69.28 51.0 37.27 23.8 62.03 +0.84
SFT Vision 62.66 83.41 81.01 63.52 87.69 68.63 50.89 36.7 38.4 63.66 +1.47
+ MergeMix 62.92 84.19 81.18 64.31 87.28 70.46 51.0 37.46 37.8 64.07 +2.88
5.3 Implementations
In this subsection, we briefly introduce the implementations on the classification task and the MLLM benchmark. The full description is in Appendix B.2. For the CIFAR100, images are resized to $224\times 224$ for ViT-based models (e.g., DeiT) and trained with AdamW (weight decay 0.05), batch size 100, for 200 or 600 epochs. We use RandomFlip and RandomCrop, plus RandAugment (cubuk2020randaugment). Learning rates are 1e-3 (DeiT-Tiny/Small, cosine schedule), 5e-4 (ViT-Small/Base), and 2e-4 (ViT-Large). For ImageNet-1K, we adopt the same settings but use a 1e-3 learning rate, batch size 1024, and 300 epochs for DeiT-Tiny/Small. For fine-grained datasets (CUB200, FGVC-Aircrafts, Stanford-Cars), we fine-tune DeiT-Small and ViT-Base for 200 epochs, batch size 16, learning rate 1e-5, using PyTorch pretrained weights (nips2019pytorch).
Following LLaVA-v1.5, we adopt Vicuna-v1.5 7B (chiang2023vicuna) as the language decoder and use a pre-trained 2-layer MLP projection to align visual and textual modalities, trained for one epoch on LCS-558K. The vision encoder is a pre-trained CLIP model that extracts image representations. During SFT, we train for one epoch on llava-v1.5-mix665k with a 2e-5 learning rate, batch size 64, and AdamW optimizer, using a 0.03 warmup ratio and cosine scheduler. Unlike LLaVA, we unfreeze the vision encoder during training. For Qwen2.5-VL-Instruction, we post training within the official checkpoints and the llava-v1.5-mix665k dataset for 0.1 epoch under similar optimization settings. The learning rates for the vision encoder, LLM decoder, and merger are set to 2e-6, 1e-5, and 1e-5, respectively.
5.4 Results of Image Classification
We did the experiments on three classification datasets on a small-scale dataset (CIFAR100), a large-scale dataset (ImageNet-1K), and a fine-grained dataset (Stanford-Cars). (1) CIFAR100: Table 1 and Table A1 shows the part and full classification results respectively. MergeMix brings the +2.15%, +2.51% gains compared with TranMix on DeiT models. Gains +2.87%, +2.88% and +4.79% on ViT models. All the results in Table 1 are trained for 200 epochs on the CIFAR100 dataset. (2) Stanford-Cars: Table 2 shows the fine-grain classification results on Stanford-Cars dataset. MergeMix achieves the 88.42% and 92.20% accuracy compared with other mixup methods. About the results of the CUB200 dataset and the FGVC-Aircrafts dataset in Table A2. (3) ImageNet-1k: Table 3 shows the results of accuracy, throughput, and flops on the ImageNet-1K dataset. It is notable that MergeMix brings a +0.27% gains and reduces -0.68G Flops compared with TransMix, and can also see other mixups with less throughput since they bring extra cost, but MergeMix has a high throughput of 1591.66 TP/s.
Figure 4: The confidence plots of mixup variants and MergeMix on the CIFAR100 dataset using DeiT-Tiny and ViT-Small. The red line indicates the expected prediction tendency.
Table 6: The calibration results of LLaVA-v1.5-7B on POPE, ScienceVQAI, GQA & SEEDI. rl denotes training with ranking loss.
Method GQA POPE SEEDI ScienceVQAI
Baseline 14.57 13.16 33.79 28.09
Training with Full Vision Tokens
SFT Vision 8.52 12.82 32.67 21.66
[…]
Training with 50% Vision Tokens
SFT Vision 18.13 12.67 34.41 24.28
[…]
Figure 5: Visualization of MergeMix on different mixing ratios, including mixed images, GradCAM of top-2 logits, and ToMe source maps.
5.5 Results of MLLM Benchmarks
We chose two mainstream MLLMs for our experiments on the VQA tasks and reasoning. Table 4 shows the results of LLaVA benchmarks, with different vision tokens for training. With the setting of full vision tokens on the training stage, our method achieves an average gain of +0.83%. When reducing the vision token to 288, our method still performs well compared with SFT. A full comparison of results in Table A10, Table A11, and Table A12. In those different settings, MergeMix achieves an average performance of 66.84%, improving over the LLaVA baseline by +1.27%. Table 5 shows the results of some VQA tasks and reasoning with Qwen2.5-VL-Instruction. MergeMix achieves an average gain of +2.88% over Qwen2.5-VL-Instruction. For the MathVista reasoning task, we reported the results without LLM-as-the-judge-eval. For the MMMU and MMMU-Pro tasks, we can achieve results on par with methods targeting reasoning improvements.
5.6 Results of Calibration
DNNs are prone to overconfidence in classification tasks. Mainfold Mixup (verma2019manifold) found that the mixup methods can effectively alleviate this problem. To this end, we compute the Expected Calibration Error (ECE (guo2017calibration)) of various mixup approaches on the CIFAR100 dataset for image classification. Also, to further analyze the calibration of MLLMs, we implement four short answer tasks, POPE, GQA, ScienceVQA, and SEED. Figure 4 shows the results of DeiT-Tiny and ViT-Small models trained for 200 epochs, showing that MergeMix obtains the best calibration of 6.7% and 9.7% in those ViT-specific mixup methods. (i.e., TransMix, SMMix, and MixPro). Table 6 shows the results of the LLaVA baseline, LLaVA with SFT, and LLaVA with our approach. SFT reduces the ECE when tuning the vision encoder, with augmentation and ranking loss, which can be better since we bring in the reward signal for the model. The more comprehensive results of CIFAR100 and MLLM benchmark we plot in Table A3 and Table A9.
5.7 Ablation Study
The ablation study mainly focuses on three things. (a) Token merge module and optimized mixing ratio, whether efficient for image classification task; (b) Exploring the ability of vision encoder and the proposed training paradigm. For the image classification scenario, Table 7 shows that compared with TopK sampling, our token merge can improve performance with +0.55%, +1.27% gains respectively, which means token merge smooths the discrete attention score. The re-scaling mixing ratio further gains +2.23%, +0.56% on the CIFAR100 dataset. For the paradigm, we validate the token merge for the LLaVA-v1.5 7B model, further explore the training with an unfrozen vision encoder, and the ranking loss. Table 8 shows that, compared with vanilla Token Merge, unfreezing the vision encoder can perform better than freezing. The augmentation and ranking loss bring more performance than only the SFT loss; (c) For further exploring the performance of hyperparameters. We also evaluated the sensitivity of hyperparameters on MergeMix, i.e., merged tokens and feature layer for better performance. Figure 6 shows the results of those hyperparameters.
Figure 6: Sensitivity analysis study of 2 hyperparameters of MergMix. Left: Different merge ratios of backbones. Right: Attention score obtained from feature layer. Those results are from training for 200 epochs.
Table 7: Ablation study of MergeMix on classification trained 200 epochs.
Module DeiT-Small ViT-Small
Vanilla 65.81 62.64
+ TopK 75.80 75.19
+ ToMe 76.45 76.46
+ Re-Scaling $\lambda$ 78.68 77.02
Table 8: Ablation study of MLLM training paradigm on LLaVA benchmark.
Method VizWiz SciVQAI MMBench
LLaVA-v1.5-7B 50.0 66.8 64.3
+ ToMe 50.45 68.86 62.8
+ SFT 48.15 69.51 66.32
[…]
6 Conclusion
This paper presents MergeMix, a unified augmentation for both image classification and MLLM alignment with token merge, also bridging the SFT and RL by building the preference pairs. Optimizing models through the mixed image and the raw image via a ranking loss. Extensive experiments demonstrate that MergeMix not only improves the performance on classic image classification tasks but also achieves a beneficial alignment and generalization on MLLM benchmarks. MergeMix provides a promising step toward a scalable, robust training paradigm for the multi-modal system.
Future Works
There remain limitations in MergeMix for MLLMs. In future work, we will address them from two directions: (1) Data level: MergeMix currently enhances only the image modality, while text remains untouched. Extending mixup to the text modality could provide more fine-grained optimization. (2) Model level: The token-merging policy is static and unlearned. Making it learnable may further improve the mixing capability.
Appendix A Declaration of LLM Usage
We use the Large Language Models (LLMs) for this paper to serve one purpose: to aid and polish the paper writing. We use the LLMs in a very limited capacity, restricted to minor editing of grammar, phrasing, and readability. We do not involve the LLMs in designing the method, developing theoretical results, and conducting experiments.
Appendix B Detail Information
B.1 Datasets
In this subsection, we will describe the datasets we chose in detail.
Image Classification.
We choose five mainstream classification datasets: (i) CIFAR100 dataset (krizhevsky2009learning) consists of 100 classes of color images with a resolution of 32 $\times$ 32 pixels, containing 50,000 training images and 10,000 test images. (ii) ImageNet-1k dataset (russakovsky2015imagenet) consists of 1,000 classes with varied image resolutions commonly cropped to 224 $\times$ 224 pixels, containing 1,281,167 training images and 50,000 test images. (iii) CUB200 dataset (wah2011caltech) contains 200 bird species, including total 11,788 images, we divided 5,994 images as training images and 5,794 images as test images, (iv) FGVC-Aircrafts dataset (maji2013fine) contains 100 aircraft model classes, including 6,667 training images and 3,333 test images, (v) Stanford-Cars dataset (jonathan2013cars) contains 196 car model classes, including 8,144 training images and 8,041 test images. All the fine-grained datasets we used set the resolution as 224 $\times$ 224 pixels for training and testing.
MLLM Benchmark.
[…]
B.2 Implementations
Classification tasks:
(i) For the CIFAR100 dataset, aiming to be suitable for training ViT-based approaches, e.g., DeiT, we resize images to $224\times 224$ and train them with the AdamW optimizer with weight decay of 0.05, batch size of 100, and total training of 200 epochs and 600 epochs. Uses RandomFlip and RandomCrop as basic augmentations, and additionally, we use RandAugment (cubuk2020randaugment). For DeiT-Tiny and DeiT-Small, we use the learning rate of 1e-3 with a dynamic cosine scheduler. For ViT-Small and ViT-Base models, we set the learning rate to 5e-4, the learning rate of ViT-Large up to 2e-4, all dynamically adjusted by a cosine scheduler. (ii) For the ImageNet-1K dataset, the dataset settings are the same as CIFAR100, but we use the learning rate as 1e-3, batch size of 1024, and a total training of 300 epochs for DeiT-Tiny and DeiT-Small with AdamW optimizer with weight decay of 0.05. (iii) For all fine-grain datasets, i.e., CUB-200 dataset, FGVC-Aircrafts dataset, and Stanford-Cars dataset, we fine-tune the DeiT-Small and ViT-Base model for 200 epochs with a batch size of 16, learning rate of 1e-5, loading the pre-trained model weight from PyTorch (nips2019pytorch).
Table A1: Top-1 accuracy (%) of mixup methods on CIFAR-100 dataset under DeiT-Tiny/Small, ViT-Small/Base/Large different model sizes. The $\alpha$ parameter of the Beta distribution follows the setting in OpenMixup (2022openmixup) setting.
Method DeiT-Tiny DeiT-Small ViT-Small ViT-Base ViT-Large
200 epochs 600 epochs 200 epochs 600 epochs 200 epochs 600 epochs 200 epochs 600 epochs 200 epochs
Vanilla 64.70 66.70 65.81 68.50 62.64 66.32 63.33 66.47 61.83
MixUp 69.47 73.06 69.98 76.35 68.67 73.57 69.66 73.90 67.90
CutMix 75.98 79.60 74.21 79.54 69.67 76.66 72.18 71.94 68.97
FMix 72.73 77.24 70.41 74.31 68.41 72.55 68.62 71.10 66.12
GridMix 71.54 76.23 68.86 74.96 70.15 68.23 66.63 68.49 63.20
ResizeMix 69.42 72.98 68.54 71.95 67.86 69.09 63.72 69.33 63.48
SaliencyMix 69.83 75.45 69.78 76.60 70.14 74.09 68.75 75.50 67.12
PuzzleMix 73.40 79.96 73.60 81.01 70.92 78.44 71.13 79.49 69.77
AutoMix 72.91 81.16 76.24 80.91 68.44 77.73 73.40 - 72.10
AdAutoMix 72.83 77.97 72.63 78.94 69.66 - 71.43 - 69.69
DeiT 74.01 79.90 75.92 79.54 72.96 77.60 72.15 76.26 69.23
TransMix 75.31 80.66 76.17 79.33 74.15 78.27 72.87 77.89 71.40
SMMix 73.84 78.62 74.09 79.84 73.50 79.65 70.87 78.18 71.38
MixPro 74.78 80.19 75.26 79.55 73.49 80.02 73.18 78.69 72.28
MergeMix 77.46 81.20 78.68 80.39 77.02 81.44 75.75 79.59 76.19
Table A2: Top-1 accuracy (%) of mixup methods on Fine-Grained datasets: CUB200, FGVC-Aircrafts, and Stanford-Cars.
Method $\alpha$ CUB200 FGVC-Aircrafts Stanford-Cars
DeiT-Small ViT-Base DeiT-Small ViT-Base DeiT-Small ViT-Base
Vanilla - 82.05 88.00 77.59 80.86 86.77 91.31
MixUp 1.0 84.31 88.75 78.52 82.18 87.73 91.36
CutMix 0.2 81.69 87.76 75.67 80.08 88.37 91.53
SmoothMix 0.2 83.87 87.02 75.31 76.72 86.39 90.88
FMix 0.2 82.64 88.68 77.08 79.33 87.18 91.36
GridMix 0.2 82.34 87.23 75.85 78.49 87.58 91.31
ResizeMix 1.0 82.15 87.61 74.59 77.62 87.45 91.59
Attentive-CutMix 2.0 82.83 87.47 75.04 76.06 87.35 90.29
SaliencyMix 0.2 82.34 87.92 77.98 79.81 87.94 91.47
PuzzleMix 1.0 84.39 88.23 78.28 81.27 88.60 91.83
[…]
DeiT 0.2 84.04 88.47 75.89 81.07 88.72 92.17
TransMix 1.0 83.34 88.10 75.73 77.77 88.38 91.66
SMMix 1.0 82.88 88.35 76.42 78.40 88.76 91.93
MixPro 1.0 82.31 86.93 75.25 75.97 88.38 91.48
MergeMix 1.0 85.40 88.40 80.92 81.97 89.42 92.20
MLLM benchmark:
Following the LLaVA-v1.5 settings, we use a pre-trained Vicuna-v1.5 7B (chiang2023vicuna) as the language decoder, which uses a pre-trained 2 $\times$ MLP as the projection for aligning the vision and text modistes, which was trained for one epoch on LCS-558K. For the vision encoder, we use a pre-trained CLIP encoder and extract the visual representation from the input images. For SFT, the learning rate was set as 2e-5, the batch size was 64, and training one epoch on llava-v1.5-mix665k dataset, uses AdamW optimizer with (0.9, 0.999) betas and epsilon of 1e-8, warmup ratio of 0.03 with a cosine scheduler. The difference from LLaVA is that we unfreeze the vision encoder during training. About Qwen2.5-VL-Instruction, we fine-tune with llava-v1.5-mix665k dataset for 0.1 epoch, uses AdamW optimizer with (0.9, 0.999) betas and epsilon of 1e-8 like LLaVA, warmup with 0.03 ratio. The learning rate of the vision encoder, LLM decoder, and merger were set to 2e-6, 1e-5, and 1e-5, respectively.
B.3 Algorithms
Algorithm 1 and Algorithm 2 show the pseudo codes for MergeMix on both image classification and preference tuning.
Algorithm 1 MergeMix for Image Classification
# Inputs: vision model $f_{\theta}(\cdot)$, training parameters $\theta$ of model, dataset $D$, mixup parameter $\alpha$; output: updated $f_{\theta}(\cdot)$
#  sample a paired mini-batch from $D$ (two images and labels)
1: for ($x_{i},y_{i}$), ($x_{j},y_{j}$) in DataLoader($D$):
#  sample mixup ratio $\lambda$ (shared with Alg. 2)
2:  $\lambda\sim\mathrm{Beta}(\alpha,\alpha)$
3:  $\mathcal{M},\hat{\lambda}=\mathcal{P}(x_{i},x_{j},\lambda)$
#  lines 3--4: MergeMix augmentation (same core as Alg. 2)
4:  $\hat{x}=\mathcal{M}\odot x\_i+(1-\mathcal{M})\odot x\_j$
[…]
#  main supervised loss on mixed sample (parallels $L_{\text{SFT}}$ in Alg. 2)
6:  $\mathcal{L}_{\text{MCE}}=\hat{\lambda}\cdot\mathcal{L}_{\text{CE}}(\textrm{logits},y_{i})+(1-\hat{\lambda})\cdot\mathcal{L}_{\text{CE}}(\textrm{logits},y_{j})$
#  optional regularizer for augmentation consistency
7:  $\mathcal{L}_{\text{CE}}=f_{\theta}(x,y)$
8:  $\mathcal{L}_{\text{Total}}=\mathcal{L}_{\text{MCE}}+\mathcal{L}_{\text{CE}}$
9:  $\mathcal{L}_{\text{Total}}$.backward()
10: optimizer.step()
11: optimizer.zero_grad()
Algorithm 2 MergeMix for MLLM Alignment
# Inputs: MLLM $\pi_{\theta}(\cdot)$, learnable parameters $\theta$ of model, dataset $D$ of $(x,q,T)$; output: updated $\pi_{\theta}(\cdot)$
#  iterate over triples
1: for $(x,q,T)$ in DataLoader($D$):
#  sample an auxiliary image $x_{i}$ (mirrors mix partner in Alg. 1)
2:  ($x_{i},$\_) = randomSample($D$)
3:  $\mathcal{M}=\mathcal{P}(x,x_{i})$
#  lines 3--4: same MergeMix augmentation core as Alg. 1
4:  $\hat{x}=\mathcal{M}\odot x+(1-\mathcal{M})\odot x_{i}$
#  winner (raw) vs. loser (mixed) outputs for the same prompt
5:  $Y_{w}=\pi_{\theta}(x,q)$
6:  $Y_{l}=\pi_{\theta}(\hat{x},q)$
#  supervised fine-tuning loss (analogous to $\mathcal{L}_{\text{CE}}$ in Alg. 1)
7:  $\mathcal{L}_{\text{SFT}}=\mathcal{L}_{\text{CE}}(Y_{\text{w}},T)$
8:  $s_{w}=\mathrm{AvgLogProb}(Y_{w},T)$
9:  $s_{l}=\mathrm{AvgLogProb}(Y_{l},T)$
#  preference loss (SimPO) that prefers winner over loser
10: $\mathcal{L}^{\text{Mix}}_{\text{SimPO}}=-\log\sigma(s_{w}-s_{l}-\lambda)$
11: $\mathcal{L}_{\text{Total}}=\mathcal{L}_{\text{SFT}}+\mathcal{L}^{\text{Mix}}_{\text{SimPO}}$
12: $\mathcal{L}_{\text{Total}}$.backward()
13: optimizer.step()
14: optimizer.zero_grad()
B.4 Future work
There are still some shortcomings in MergeMix for MLLMs. In future work, we will explore this from two levels: (1) From the data perspective: MergeMix focuses on the enhancement of the image modality during training, while text inputs still remain raw. How to extend mixup to the text modality in MLLM tasks needs to be solved, as this can provide more fine-grained optimization guidelines. (2) From the model perspective: The token merging is still static and unlearned. Improving the merging strategy to make it learnable via metrics or backpropagation could enhance the token merging ability for mixing.
Appendix C Image Classification
C.1 Extensive Results
Figure A1: Robustness against image occlusion classification results with different occlusion ratios for different mixup methods based on DeiT-Tiny (left) and DeiT-Small (right) on CIFAR100 and ImageNet-1K datasets.
Table A1 shows the full results of 200 epochs and 600 epochs training on the CIFAR100 dataset using different ViT models. Table A2 shows the three fine-grain datasets, i.e., CUB200, FGVC-Aircrafts, and Stanford-Cars. It is easily found that MergeMix can achieve the SOTA on lots of models. Also, the speed of overfitting is a significant improvement over other methods, which means the token merge can gather useful information and reduce some redundant tokens.
Table A3: The calibration results of ViT-based mixup methods on CIFAR-100 & ImageNet-1K, with training 200 and 300 epochs respectively. tome denotes inference with token merging.
[…]
DeiT-Tiny CIFAR100 200 8.64 11.42 14.52 12.24 11.80 10.68 7.08 6.73
ViT-Small CIFAR100 200 13.89 14.43 14.22 15.76 12.41 10.65 9.69 10.14
ViT-Large CIFAR100 200 4.76 7.36 16.44 6.71 7.70 6.22 4.65 4.70
DeiT-Tiny CIFAR100 600 6.44 6.01 7.56 8.53 8.35 6.38 5.45 5.46
ViT-Small CIFAR100 600 9.01 6.45 9.12 10.22 9.55 7.42 5.50 6.04
DeiT-Small ImageNet-1K 300 5.66 4.19 7.57 8.24 6.53 5.17 6.00 4.52
Table A4: The Top-1 accuracy of DeiT-Tiny trained by various Mixup approaches on the CIFAR100 dataset with different occlusion ratios. tome denotes inference with token merging.
Method DeiT-Tiny Trained 200 epochs
0% 10% 20% 30% 40% 50% 60% 70% 80% 90%
Vanilla 66.68 64.54 62.57 60.20 56.96 51.41 40.32 27.75 13.25 4.99
MixUp 73.06 70.52 68.45 65.39 61.27 53.55 42.80 29.77 16.01 5.01
CutMix 79.58 78.64 77.10 75.83 72.96 70.25 64.40 54.39 35.82 14.08
FMix 77.14 76.01 74.62 73.33 71.27 67.72 63.17 56.18 42.12 17.32
GridMix 76.13 73.79 71.94 69.36 66.36 62.34 56.02 47.30 32.52 14.35
ResizeMix 72.93 71.82 70.53 69.67 67.77 65.22 59.87 50.26 31.26 9.72
SaliencyMix 75.41 74.63 73.57 72.14 69.47 65.05 58.08 44.82 24.14 7.03
Attentive-CutMix 80.27 79.13 77.94 76.98 75.45 71.01 57.75 33.10 12.03 4.02
PuzzleMix 79.97 78.18 77.36 76.04 73.74 70.74 65.83 56.75 40.43 19.41
AutoMix 81.12 78.37 78.40 78.16 77.65 77.09 74.63 71.64 67.52 55.61
TransMix 80.66 79.77 75.33 73.60 70.98 69.95 63.20 40.02 25.28 9.77
DeiT 79.80 78.20 77.29 75.83 74.59 72.31 69.05 60.18 38.55 12.53
SMMix 78.62 77.78 76.51 74.78 71.25 64.17 47.28 24.72 7.12 2.19
MixPro 80.14 79.61 78.78 77.06 74.81 71.18 64.25 50.75 26.80 5.19
MergeMix 80.88 75.19 69.71 68.72 69.33 70.17 69.36 68.20 63.49 51.72
[…]
Table A5: The Top-1 accuracy of ViT-Small trained by various Mixup approaches on the CIFAR100 dataset with different occlusion ratios. tome denotes inference with token merging.
Method ViT-Samll Trained 200 epochs
0% 10% 20% 30% 40% 50% 60% 70% 80% 90%
Vanilla 66.24 62.51 60.37 58.59 56.68 53.61 49.70 39.70 22.60 7.55
MixUp 73.67 71.84 70.80 69.07 66.75 63.52 58.21 48.68 28.72 8.01
CutMix 76.13 73.77 72.93 71.92 70.32 68.32 65.48 59.30 43.11 17.04
FMix 71.27 68.52 66.49 65.78 65.00 62.31 57.80 48.98 30.09 6.19
GridMix 67.99 66.48 65.60 64.21 62.54 60.40 56.98 50.02 34.62 9.74
ResizeMix 66.69 66.90 62.80 58.15 47.00 36.15 36.49 33.07 10.97 2.06
SaliencyMix 73.50 72.12 71.88 71.21 69.97 67.62 63.03 53.12 33.27 11.06
Attentive-CutMix 78.26 72.49 67.52 63.83 60.76 54.93 38.05 27.75 32.20 31.66
PuzzleMix 78.01 76.42 75.54 74.61 73.99 71.46 68.23 63.24 51.51 26.25
AutoMix 77.52 76.74 75.14 72.95 69.71 66.76 62.59 59.00 49.01 27.45
TransMix 78.37 76.29 75.86 75.58 74.79 73.19 70.81 68.18 62.29 51.09
DeiT 77.50 75.99 75.80 74.57 73.80 71.73 67.23 58.40 43.49 19.94
SMMix 79.32 77.84 76.60 75.41 73.00 68.10 58.89 45.07 23.25 4.59
MixPro 79.91 78.37 75.36 72.08 68.37 57.48 37.08 21.62 11.87 3.69
MergeMix 81.29 75.38 73.41 69.36 61.02 51.73 52.52 62.14 64.83 57.21
[…]
Table A6: The Top-1 accuracy of DeiT-Small trained by various Mixup approaches on the ImageNet-1K dataset with different occlusion ratios. tome denotes inference with token merging.
Method DeiT-Samll Trained 300 epochs
0% 10% 20% 30% 40% 50% 60% 70% 80% 90%
Vanilla 75.46 74.03 72.85 71.36 69.91 67.94 64.71 60.42 51.65 34.08
MixUp 78.74 78.17 77.11 76.21 74.31 71.85 68.13 62.40 51.45 33.08
CutMix 80.16 79.23 78.20 76.87 75.21 72.92 69.24 64.63 55.95 39.69
TransMix 80.36 79.47 78.24 77.00 75.40 73.31 70.17 65.78 57.79 42.44
DeiT 80.27 79.03 77.92 76.39 74.65 72.25 68.31 63.53 54.57 38.81
SMMix 79.32 78.56 77.70 76.17 73.76 70.25 64.95 58.94 48.69 34.37
MixPro 79.25 78.83 78.01 77.24 76.02 74.26 71.48 67.44 58.49 39.47
MergeMix 80.70 80.38 79.95 78.97 78.08 76.21 74.08 70.71 65.57 51.98
[…]
C.2 Robustness Experiments of Mixup Augmentations
Calibration of mixup augmentations
Table A3 shows the results of calibration on seven mixup augmentations. We evaluated with two public datasets by DeiT-Tiny, ViT-Small, and ViT-Large on the CIFAR100 dataset, by models trained with 200 epochs, DeiT-Tiny and ViT-Small trained with 600 epochs, and DeiT-Small on the ImageNet-1K dataset, trained with 300 epochs.
Results of occlusion robustness
The full results of occlusion robustness classification on MergeMix and other mixup methods. Figure A1 shows the curve of MergeMix and other mixup methods on the CIFAR100 dataset and ImageNet-1K dataset. Table A4 and Table A5 show the accuracy results on the CIFAR100 dataset by vanilla and 14 different mixup approaches. Table A6 shows the results of 8 different methods on the ImageNet-1K dataset.
Table A7: Classification results of CAFormer small (CAFormer-S12) with different mxiup augmentations, training 200 epochs with 100 batch size on CIFAR100 dataset.
Model Dataset Epochs Vanila MixUp CutMix DeiT TransMix MergeMix
CAFormer-S12 CIFAR100 200 74.95 81.64 84.69 83.60 83.70 84.30
C.3 Further results on different backbone and data modality.
Table A8: Supervised fine-tuning on HuBERT-Base model on ESC-50 and UbranSound8K datasets.
HuBERT-Base ESC-50 UbranSound8K
Vanilla 75.12±1.07 84.14±0.45
MixUp 75.86±0.83 85.02±0.26
TransMix 76.27±1.14 85.33±0.57
MergeMix 76.51±0.95 85.69±0.42
For further exploring the effectiveness of MergeMix, we applied our approach on a hyper-model MetaFormer (CAFormer) (yu2022metaformer) and two audio datasets for classification by HuBERT-Base (hsu2021hubert). ESC-50 (piczak2015esc) consisted of 50 classes, with 1,200 training samples and 400 validation samples, and a maximum duration of 3 seconds. UrbanSound8k (salamon2014urban8k) is a classification dataset consisting of 10 classes, with a maximum duration of 4 seconds, containing 7,079 training samples and 816 validation samples. Following the USB experimental setup, we fine-tuned the model for 100 epochs using the AdamW optimizer. The base learning rate was set to 1e-4, texttt5e-4, with a batch size of 32 and a weight decay coefficient of 5e-4. Both the compared shuffling methods and our proposed MergeMix can be directly transferred to audio data (treated as one-dimensional sequences). Table A8 below shows our reproduced comparison results on two audio datasets. Compared to the MixUp and TransMix baselines, MergeMix achieves significant performance improvements over multiple shuffling baseline models. For CAFormer-S12, we train for 200 epochs on the CIFAR100 dataset with the same settings as DeiT-Tiny. Table A7 shows that MergeMix obtained a second-best result, since CAFormer only has with Attention module of 2 stages (8 layers).
Appendix D Visual Understanding
D.1 Calibration Results of MLLM
Table A9: The calibration results of LLaVA on POPE, ScienceVQAI, GQA & SEEDI.
Method GQA POPE SEEDI ScienceVQAI
Baseline 14.57 13.16 33.79 28.09
Training with Full Vision Tokens
SFT Vision 8.52 12.82 32.67 21.66
[…]
Training with 50% Vision Tokens
SFT Vision 18.13 12.67 34.41 24.28
[…]
Training with 25% Vision Tokens
SFT Vision 13.32 12.51 34.97 18.86
[…]
To explore MLLM calibration, we selected 4 tasks with short responses across different token-reduction settings. We evaluate three scenarios: using full vision tokens, 50% tokens, and 25% tokens, and compare various data augmentation strategies (MixUp, CutMix, ResizeMix, MergeMix) trained with ranking loss (denoted as rl). Table A9 shows that with unfreezing the vision encoder, the ECE can be better than freezing. With the vision tokens reduced in training. Overall, token reduction leads to a moderate drop in calibration accuracy, but effective augmentation strategies significantly mitigate this degradation. In the full-token setting, CutMixrl achieves the lowest GQA calibration error (6.09), while ResizeMixrl shows the best SEED result (36.08). When reducing tokens to 50%, CutMixrl and MergeMixrl remain competitive, maintaining strong calibration across tasks despite reduced visual information. Even with 25% tokens, CutMixrl continues to yield relatively balanced performance, indicating that appropriate augmentation enhances robustness under token compression. These results suggest that MLLM calibration remains an open question, especially in environments that require a reliable answer.
D.2 Relationship between Mixing Ratios and Rewards
Figure A2: LLM-as-the-Judge results for different mixing ratios. Higher mixing ratios consistently yield higher reward scores across strong MLLMs. MergeMix aligns closely with these models, showing that the mixing ratio provides a meaningful preference signal.
To understand how the proposed ranking loss on synthetic mixed pairs approximates human preference learning, we conduct a LLM-as-the-judge-eval evaluation in which strong MLLMs score the mixed pairs generated with different mixing ratios $\lambda$. Specifically, we query a diverse set of frontier MLLMs, including Grok-3, Doubao-Seed-1.6 (guo2025seed1), Doubao-Seed-1.6-CoT, Qwen3-VL-Plus-32B (yang2025qwen3), Qwen2.5-VL-72B (bai2025qwen2.5), Gemini-2.5 (team2023gemini), and a human-expert baseline, to assign reward scores to each mixed pair. As shown in Figure A2, we found that the reward consistently increases with larger mixing ratios, and this trend holds across nearly all evaluators. The strong monotonic correlation suggests that $\lambda$ provides a reliable and well-behaved control signal for modeling preference strength. This observation validates our use of $\lambda$ as an interpretable proxy for preference supervision and highlights its potential as a lightweight alternative to explicit human-annotated reward signals.
D.3 Results of Different Vision Token Ratios on Inference
In this subsection, we validate the different ratios of vision tokens on the LLaVA benchmark. Table A10, Table A11 and Table A12 show the fully results with full, 75%, 50% and 25% ratios on inference, respectively. Those results give a full comparison of the influence on the vision tokens. Significantly shown in Table A10, MergMix always brings gains in different merge ratios, from +0.83 to +0.06. Other methods, since they are highly random, cause performance instability.
From the results shown in Table A11, when the training stage uses the token merge, it can achieve an average gain of 66.84%, which improves 0.43% over training and inference without token merge training and inference, with an improvement of +1.27% than the average performance of the original LLaVA model. Figure A3 shows the average accuracy on the LLaVA benchmark for the LLaVA-v1.5-7B model trained and evaluated under different vision token-merging ratios. The results demonstrate that MergeMix maintains strong performance across a wide range of settings, outperforming or matching other baselines.
Table A10: Different token merge ratios of inference comparison results with augmentations. AVG: The average of the nine benchmarks for comprehensive comparison, except for MME.
LLaVA-7b Train Image Question Answering Benchmarks AVG Gain
[…]
Inference Full Vision Token
Vanilla - 78.5 62.0 50.0 66.8 58.2 1510.7 64.3 58.3 85.87 66.19 65.57 -
SFT Vision 100% 79.32 62.98 47.45 70.05 57.17 1490.88 66.26 60.05 86.18 67.32 66.31 +0.74
+ MixUp 100% 79.27 62.58 44.95 69.41 57.39 1483.20 65.72 58.24 86.27 66.73 65.62 +0.05
+ CutMix 100% 79.18 62.40 45.04 70.60 57.06 1452.31 66.32 58.24 86.47 67.22 65.84 +0.27
+ ResizeMix 100% 77.78 61.66 44.43 68.91 55.11 1436.09 63.91 55.41 86.01 63.91 64.13 -1.44
+ MergeMix 100% 79.24 62.44 47.69 69.86 57.56 1479.97 66.58 60.65 86.10 67.47 66.40 +0.83
Inference 75% Vision Token
Vanilla - 77.24 59.65 50.86 68.42 55.66 1460.88 63.05 57.9 85.60 65.21 64.84 -
SFT Vision 100% 77.62 59.73 46.57 70.15 54.83 1454.99 65.03 59.36 85.60 65.90 64.98 +0.14
+ MixUp 100% 77.66 59.37 44.15 69.91 56.18 1457.98 64.77 57.73 85.15 65.28 64.47 -0.37
+ CutMix 100% 77.67 59.21 44.25 69.66 54.84 1400.49 65.03 57.98 85.91 65.66 64.47 -1.09
+ ResizeMix 100% 76.45 59.65 43.31 68.82 53.14 1426.75 63.91 55.15 85.29 63.04 63.20 -1.64
+ MergeMix 100% 77.71 59.32 47.46 70.70 54.85 1440.50 65.37 58.93 85.04 65.98 65.04 +0.20
Inference 50% Vision Token
Vanilla - 76.65 59.33 50.45 68.86 55.33 1452.18 62.8 56.87 86.53 64.09 64.55 -
SFT Vision 100% 77.07 59.05 46.19 70.35 54.40 1436.79 64.60 58.93 85.62 65.12 64.59 +0.04
+ MixUp 100% 77.11 58.96 44.39 69.36 54.41 1422.28 64.34 57.90 86.16 64.29 64.10 -0.45
+ CutMix 100% 77.15 58.66 44.00 69.31 54.78 1428.21 64.94 59.02 86.42 64.80 64.34 -0.21
+ ResizeMix 100% 75.90 59.46 42.99 69.31 52.83 1444.13 63.23 54.20 85.77 62.07 62.86 -1.69
+ MergeMix 100% 77.13 59.02 47.46 70.55 54.54 1461.63 65.03 58.84 85.70 65.25 64.84 +0.29
Inference 25% Vision Token
Vanilla - 74.63 58.76 52.71 68.67 55.32 1398.24 60.65 54.03 86.54 62.05 63.71 -
SFT Vision 100% 75.02 58.43 46.36 69.06 52.32 1376.88 62.37 55.84 85.64 63.63 63.19 -0.52
+ MixUp 100% 75.45 58.63 44.82 68.02 52.13 1384.31 62.28 60.65 85.54 62.97 63.39 -0.32
+ CutMix 100% 75.39 58.61 43.85 68.67 52.69 1330.99 62.71 56.01 86.30 63.02 63.03 -0.68
+ ResizeMix 100% 74.08 59.02 43.40 68.96 51.40 1377.96 61.08 53.09 86.53 59.85 61.93 -1.78
+ MergeMix 100% 75.50 58.60 48.14 69.86 52.01 1439.44 63.05 57.56 85.53 63.72 63.77 +0.06
Table A11: Different token merge ratios of inference comparison results with augmentations. AVG: The average of the nine benchmarks for comprehensive comparison, except for MME.
LLaVA-7b Train Image Question Answering Benchmarks AVG Gain
[…]
Inference Full Vision Token
Vanilla - 78.5 62.0 50.0 66.8 58.2 1510.7 64.3 58.3 85.87 66.19 65.57 -
SFT Vision 50% 78.6 62.47 48.15 69.51 56.41 1486.24 66.32 57.98 87.37 66.75 65.95 +0.38
+ MixUp 50% 78.51 62.07 51.1 68.47 56.54 1459.06 65.63 59.53 86.86 66.06 66.08 +0.51
+ CutMix 50% 78.58 62.39 50.53 70.2 55.95 1414.72 66.92 59.53 86.56 66.2 66.31 +0.74
+ ResizeMix 50% 76.39 61.05 45.48 68.07 54.60 1447.35 63.31 51.97 86.57 62.54 63.33 -2.24
+ MergeMix 50% 78.61 62.18 52.14 69.61 56.85 1453.97 66.58 59.02 86.47 66.63 66.45 +0.88
Inference 75% Vision Token
Vanilla - 77.24 59.65 50.86 68.42 55.66 1460.88 63.05 57.90 85.60 65.21 64.84 -
SFT Vision 50% 78.75 62.82 48.02 70.65 56.33 1486.24 66.40 59.02 86.93 67.17 66.23 +1.39
+ MixUp 50% 78.87 62.32 51.01 69.11 56.62 1480.04 65.63 59.53 86.86 66.06 66.22 +1.38
+ CutMix 50% 78.73 62.42 49.85 70.50 56.12 1418.07 67.61 59.87 85.96 66.44 66.39 +1.55
+ ResizeMix 50% 76.79 61.12 44.85 68.37 54.24 1475.27 64.26 52.66 85.67 63.30 63.47 -1.37
+ MergeMix 50% 78.81 62.50 52.31 69.56 56.51 1455.81 66.66 59.10 85.76 67.12 66.48 +1.64
Inference 50% Vision Token
Vanilla - 76.65 59.33 50.45 68.86 55.33 1452.18 62.80 56.87 86.53 64.09 64.55 -
SFT Vision 50% 78.49 63.39 46.69 70.25 55.68 1468.38 66.83 57.76 86.47 66.48 65.78 +1.23
+ MixUp 50% 78.54 61.91 51.01 69.61 55.76 1468.14 65.63 59.87 86.51 66.39 66.14 +1.59
+ CutMix 50% 78.50 62.18 48.79 70.50 55.83 1431.32 67.18 59.02 86.00 66.27 66.03 +1.48
+ ResizeMix 50% 76.55 60.79 44.20 68.32 54.21 1470.22 63.31 52.66 86.23 62.73 63.22 -1.33
+ MergeMix 50% 78.51 62.09 51.01 70.10 56.03 1464.00 66.75 59.45 86.05 66.39 66.26 +1.71
Inference 25% Vision Token
Vanilla - 74.63 58.76 52.71 68.67 55.32 1398.24 60.65 54.03 86.54 62.05 63.71 -
SFT Vision 50% 77.30 61.77 46.33 70.55 54.19 1411.01 65.72 58.07 86.34 65.46 65.08 +1.37
+ MixUp 50% 77.28 61.56 48.77 69.56 54.10 1419.71 66.32 56.27 86.57 65.32 65.08 +1.37
+ CutMix 50% 77.20 61.52 49.00 71.24 53.96 1372.66 66.49 58.76 86.38 65.24 65.53 +1.82
+ ResizeMix 50% 75.06 59.88 43.12 67.13 52.34 1445.26 61.51 51.63 86.03 61.59 62.03 -1.68
+ MergeMix 50% 77.20 61.81 51.66 70.35 54.47 1401.58 66.83 59.10 85.92 65.44 66.84 +3.13
Table A12: Different token merge ratios of inference comparison results with augmentations. AVG: The average of the nine benchmarks for comprehensive comparison, except for MME.
LLaVA-7b Train Image Question Answering Benchmarks AVG Gain
[…]
Inference Full Vision Token
Vanilla - 78.5 62.0 50.0 66.8 58.2 1510.7 64.3 58.3 85.87 66.19 65.57 -
SFT Vision 25% 77.92 62.01 50.30 69.11 55.03 1420.69 64.17 56.09 86.82 65.04 65.17 -0.40
+ MixUp 25% 77.89 62.01 52.53 70.20 55.90 1444.45 64.29 57.73 86.94 65.77 65.92 +0.35
+ CutMix 25% 77.92 61.57 51.02 69.21 55.43 1408.18 64.23 57.13 86.09 65.19 65.31 -0.26
+ ResizeMix 25% 75.38 59.77 41.38 66.78 53.45 1430.64 62.37 51.54 85.26 61.26 61.91 -3.66
+ MergeMix 25% 77.86 61.54 50.50 69.56 55.40 1458.49 64.86 57.98 87.22 65.10 65.56 -0.01
Inference 75% Vision Token
Vanilla - 77.24 59.65 50.86 68.42 55.66 1460.88 63.05 57.9 85.60 65.21 64.84 -
SFT Vision 25% 78.09 62.11 49.67 69.86 55.03 1423.27 64.94 57.47 85.86 65.77 65.42 +0.58
+ MixUp 25% 77.97 61.23 53.13 69.46 56.17 1466.33 66.06 58.67 86.14 65.67 66.06 +1.22
+ CutMix 25% 78.11 61.85 50.46 69.71 56.08 1420.67 67.52 59.19 85.14 65.42 65.94 +1.1
+ ResizeMix 25% 76.01 59.95 41.94 66.63 53.75 1459.84 63.48 52.92 84.99 62.08 62.42 -2.42
+ MergeMix 25% 78.07 61.42 50.17 70.2 55.96 1483.82 66.58 59.45 86.18 65.49 65.95 +1.11
Inference 50% Vision Token
Vanilla - 76.65 59.33 50.45 68.86 55.33 1452.18 62.8 56.87 86.53 64.09 64.55 -
SFT Vision 25% 77.99 61.77 48.79 70.10 54.91 1443.02 64.69 57.38 86.21 65.41 65.25 +0.7
+ MixUp 25% 77.89 61.68 50.85 70.05 56.0 1448.62 66.92 58.07 86.44 65.54 65.94 +1.39
+ CutMix 25% 77.32 61.62 48.86 69.96 55.97 1428.02 67.01 59.02 85.89 65.57 65.69 +1.14
+ ResizeMix 25% 75.78 59.95 40.82 67.33 53.44 1456.66 63.4 53.09 85.57 61.65 62.34 -2.21
+ MergeMix 25% 77.91 61.56 48.86 70.5 56.0 1477.47 66.15 58.76 86.41 65.19 65.71 +1.16
Inference 25% Vision Token
Vanilla - 74.63 58.76 52.71 68.67 55.32 1398.24 60.65 54.03 86.54 62.05 63.71 -
SFT Vision 25% 76.97 61.43 49.79 70.0 53.56 1405.86 64.94 56.7 86.81 64.51 64.97 +1.26
+ MixUp 25% 76.84 61.15 49.31 69.66 53.98 1409.16 65.8 57.81 86.72 64.49 65.08 +1.37
+ CutMix 25% 76.89 61.3 48.21 69.32 53.99 1370.27 66.32 57.98 86.42 64.21 64.96 +1.25
+ ResizeMix 25% 75.01 59.72 40.82 66.78 51.66 1418.27 62.45 52.66 85.87 60.54 61.72 -1.99
+ MergeMix 25% 76.91 61.0 48.79 70.2 54.51 1441.07 66.06 58.24 86.57 64.2 65.16 +1.45
Figure A3: The plots of the LLaVA-v1.5-7B model under different inference time merge ratios for various methods (Baseline, SFT, MixUp, CutMix, ResizeMix, and MergeMix) demonstrate that MergeMix maintains robust performance across a wide range of configurations.
D.4 Results of Different Ranking Loss
To understand the effectiveness of ranking loss on preference tuning, we conducted an ablation study for $\mathcal{L}^{\text{Mix}}_{\text{SimPo}}$. Table A13 shows that compared with vanilla SimPO, our approach can bring more improvement.
Table A13: Ablation study of different ranking loss on LLaVA-v1.5-7B.
Models SciVQAI TextVQA VizWiz MMBench Avg. Gains
LLaVA-v1.5-7B 66.8 58.2 50.0 64.3 59.82 -
mDOP (wang2024mdpo) 67.53 57.90 50.04 64.60 60.02 +0.20
Re-Align (xing2025re) 68.10 58.55 50.06 64.69 60.35 +0.53
vanilla SimPO 69.86 56.62 49.26 66.24 60.49 +0.67
+ MergeMix 69.86 57.56 47.69 66.58 61.29 +1.47
Appendix E Efficiency
E.1 Results of Different Vision Token Ratios on Inference
To further validate the inference efficiency gains achieved by Token Merge in MergeMix, we conducted experiments on both image classification models and multi-modal large models. As shown in Table A14, increasing the merging ratio $r$ of the ViT-L model significantly reduces FLOPs (from 59.57G to 34.93G) while throughput improves from 122.83 to 201.07 (+63.7%). Moreover, the additional overhead of Token Merge itself is extremely low (only 0.97 ms at $r$ = 0.75), far below the computational cost of pre-layering. This demonstrates that Token Merge can efficiently compress visual tokens while maintaining negligible overhead. For multi-modal inference, Table A15 shows the effectiveness of Token Merge on LLaVA-v1.5-7B. As $r$ increases, the model’s overall throughput improves, while TTFT decreases significantly. The optimal trade-off is achieved at $r$ = 0.5, boosting throughput from 45.96 to 47.28 and reducing TTFT from 86.44 ms to 66.74 ms. When $r$ = 0.75, TTFT further decreases to 62.25 ms. This demonstrates that Token Merge effectively accelerates multi-modal inference without compromising image-text alignment quality. These results collectively show that Token Merge can efficiently reduce the number of visual tokens in MergeMix, thereby lowering FLOPs and inference latency in both visual models and MLLMs. This validates the necessity and advantages of introducing Token Merge into our methodology.
Table A14: Results of throughput, FLOPs (G), overhead of pre-layer and ToMe by ViT-L model with different merged ratios, evaluated on an Nvidia A100 GPU.
Ratios Throughput $\uparrow$ FLOPs (G) Overhead
Layer ToMe
Baseline 122.83 59.57 10.19 -
$r$ = 0.1 130.45 56.29 9.82 3.72
$r$ = 0.25 145.98 49.37 8.23 3.14
$r$ = 0.5 167.54 42.65 5.52 1.79
$r$ = 0.75 201.07 34.93 3.23 0.97
Table A15: Results of throughput and Time-To-First-Token (TTFT) in ms on LLaVA-v1.5-7B with different merged ratios $r$, which is evaluated on a Nvidia A100 GPU.
Ratios Throughput $\uparrow$ TTFT (ms)
LLaVA-V1.5 45.96 86.44
$r$ = 0.1 43.94 89.32
$r$ = 0.25 46.99 83.65
$r$ = 0.5 47.28 66.74
$r$ = 0.75 46.92 62.25
Appendix F Visualization and Case Study
In this section, we provide a visualization of the case study of augmentation samples with corresponding performance reward scores and an extensive visualization of mixing samples and ToMe source maps. Firstly, we provide a case of different degrees of augmentation in Figure A4. Then, we plot some visualizations of token merge with different merge ratios, mixed samples with different $\lambda$ in Figure A5 and Figure A6. For every three rows in Figure A5 and Figure A6, the first and second rows of source maps could directly capture the important regions of the raw images, where a large merge ratio enables better grouping of the similar regions. Based on the source maps, when the mixing ratio $\lambda$ grows from small to large, MergeMix could keep the most distinguishable tokens of the first image while gradually expanding more tokens from less important regions, which enables MergeMix to generate reliable mixing samples with no more computational costs. Moreover, we provide GradCAM (selvaraju2017gradcam) visualization of the top-1 and top-2 classes with the mixed samples of MergeMix on ImageNet-1K, as shown in Figure A7 and Figure A6.
Figure A4: The visualization of the visual question answers with different mixing ratios by LLaVA-v1.5-7B model. Note that the blue texts denote the core question and the corresponding correct answers, while the green texts denote the wrong answer to the question. The raw image denotes without any augmentations, and other images denote with different mixing ratios $\lambda$. Ground-truth Answer denotes the raw labels for this case. With the mixing degree improving, the answer comes out more wrong or unrelated to the question, as shown in green color.
Figure A5: Visualization of mixed samples with source maps of ToMe with different mixing ratios $\lambda$ and various merge ratios on ImageNet-1K.
Figure A6: Visualization of mixed samples with source maps of ToMe with different mixing ratios $\lambda$ and various merge ratios on ImageNet-1K.
Figure A7: Visualization of mixed samples and corresponding GradCAM (selvaraju2017gradcam) of the top-1/2 class with MergeMix on ImageNet-1K.
Figure A8: Visualization of mixed samples and corresponding GradCAM (selvaraju2017gradcam) of the top-1/2 class with MergeMix on ImageNet-1K.